OpenStudy (anonymous):

What is a quadratic formula? What is it used for? Please provide an example????

OpenStudy (anonymous):

Anyone help me. I have been out due to the flu. I have a question to answer due today, I haven't got to read that far yet. The question is What is a quadratic formula? What is it used for and provide an example? Any help?

OpenStudy (anonymous):

Still stu

OpenStudy (anonymous):

ck

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm In general sense, it solves quadratic equations quickly and properly and usually ends up giving you two answers, depending on the equation, both or just one answer will be correct.

OpenStudy (mathstudent55):

The quadratic formula is a formula that allows you to solve any quadratic equation. If you have a quadratic equation in the form \[ax ^{2} + bx + c = 0\] where a, b, and c are numbers and x is the variable, the the quadratic formula, \[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] will give you the solution. Example: Solve 2x^2 + 5x - 3 = 0 using the quadratic formula. Here you have a = 2, b = 5, c = -3 Just plug in those values of a, b, and c in the quadratic formula to get the value of x: \[x = \frac{ -5 \pm \sqrt{(5)^{2} - 4(2)(-3)} }{ 2(2) }\] \[x = \frac{ -5 \pm \sqrt{25 + 24} }{ 4 }\] \[x = \frac{ -5 \pm \sqrt{49} }{ 4 } = \frac{ -5 \pm 7 }{ 4 }\] x = 2/4 or x = -12/4 x = 1/2 or x = -3

OpenStudy (anonymous):

Thank You. These area of algebra is becoming really difficult for me to understand. I appreciate the help.

OpenStudy (mathstudent55):

You're welcome.